does this system have a unique solution? yes or no?
20x+5y=33
-32x+8y=47

Found 2 solutions by nyc_function, jsmallt9:
Answer by nyc_function(2741)

After working out the question on paper, I got x = 29/320 and y = 499/80.


Answer by jsmallt9(3758)

Since you posted this problem under "Matrices and Determinants" I assume you know what they are.

The fast way to determine whether or not there is a unique solution is to find the value of the following determinant:
|  20   5  |
| -32   8  |

The value of this determinant is 20*8 - (-32)*5 = 160 + 160 = 320. Since this determinant is non-zero there is a unique solution. (If this determinant had worked out to be zero, then there would not be a unique solution.)